# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@ 2. Case 2
(A.19) $\Omega_{n}=B_{R_{n}}(0)-\bigl{(}B_{\frac{1}{n}}(\theta_{a})\cup B_{\frac{1}{n}}(\theta_{b})\bigr{)}$
with $n>\frac{1}{\delta}$ and $R_{n}>n$ large enough such that $\int_{B_{R_{n}}^{c}(0)}\rho\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$ and $\int_{B_{R_{n}}^{c}(0)}{\rho^{*}}\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$. Without loss of generality, we assume $\int_{\Omega_{n}}{\rho^{*}}\mathrm{d}\theta>\int_{\Omega_{n}}\rho\mathrm{d}\theta$, then we define
(A.20) $\Omega_{n}(r)=B_{R_{n}}(0)-\bigl{(}B_{r}(\theta_{a})\cup B_{1/n}(\theta_{b})\bigr{)}$
@@
(A.21) $\int_{{\Omega}_{n}(r)}{\rho^{*}}\mathrm{d}\theta$ $=1-\int_{B_{R_{n}}(0)^{c}}{\rho^{*}}\mathrm{d}\theta-\int_{B_{1/n}(\theta_{b})}{\rho^{*}}\mathrm{d}\theta$
$\leq 1-\int_{B_{1/n}(\theta_{b})}(\rho-\epsilon)\mathrm{d}\theta$
$\leq 1-\int_{B_{1/n}(\theta_{b})}\rho\mathrm{d}\theta-\epsilon|B_{1/n}(0)|$
@@
Therefore, by continuity, there exists some $r_{n}\in(0,1/n)$, such that $\int_{{\Omega}_{n}(r_{n})}{\rho^{*}}\mathrm{d}\theta=\int_{{\Omega}_{n}(r_{n})}\rho\mathrm{d}\theta$. Using ${\Omega}_{n}(r_{n})$ to replace $\Omega_{n}$ completes the proof. We can also make $\Omega_{n}$ monotonically increasing by using the union $\cup_{k=1}^{n}\Omega_{k}$ to replace $\Omega_{n}$. ∎
@@
(B.1) $f(y)\geq f(x)+(y-x)f^{\prime}(x),\quad\forall\ y>0,\ x>0.$
@@
$\bigl{(}f^{\prime}(y)-f^{\prime}(x)\bigr{)}(y-x)\geq(y-x)f^{\prime}(y)-f(y)+f(x)\geq 0,\quad\forall\ y>0,\ x>0.$
Then for $x>0$, we have
(B.2) $\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(\frac{1}{\mu}-\frac{1}{x})\geq$ $(\frac{1}{\mu}-\frac{1}{x})f^{\prime}(\frac{1}{\mu})-f(\frac{1}{\mu})+f(\frac{1}{x})\geq 0,$
@@
(B.3) $=$ $\mu(x-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}x\bigl{(}f(\frac{1}{\mu})-f(\frac{1}{x})\bigr{)}.$
Taking $x={\rho^{*}}/\rho$ in (B.3) leads to
$\Bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr{)}\Bigl{(}\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr{)}\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})),$
which leads to the desired inequality. This completes the proof. ∎
Proof of Lemma 5.5.
Since $f\in C^{2}$ and $f^{\prime\prime}(1)>0$, there exists $0<\delta_{f}<\frac{1}{4}$, such that for any $x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})$, we have
(B.4) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x).$
For any $x,y$ satisfying $y>x>0$ and $\{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$, we have
(B.5) $f^{\prime}(\frac{1}{x})-f^{\prime}(\frac{1}{y})=\int_{1/y}^{1/x}f^{\prime\prime}(\xi){\rm d}\xi\geq\frac{f^{\prime\prime}(1)}{2}\int_{(\frac{1}{y},\frac{1}{x})\cap(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})}{\rm d}\xi$
@@
$\geq$ $\alpha_{f}^{\prime}\frac{y-x}{y+x},$
where
$\alpha_{f}^{\prime}=\frac{f^{\prime\prime}(1)\delta_{f}}{2(1+2\delta_{f})(1+\delta_{f})}>0.$
These cases cover all situations with $y>x>0,\ \{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$.
Then we consider (5.9) with $\mu\in[1,1+\delta_{f}]$. When $0<x<\mu$, by using (B.5), we have
(B.6) $\begin{split}\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}\frac{\frac{1}{x}-\frac{1}{\mu}}{\frac{1}{x}+\frac{1}{\mu}}(\mu^{2}-x^{2})=\alpha_{f}^{\prime}(\mu-x)^{2}.\end{split}$
When $\mu<x$, by using (B.5), we have
(B.7) $\begin{split}\bigl{(}f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr{)}(x^{2}-\mu^{2})\geq\alpha_{f}^{\prime}\frac{\frac{1}{\mu}-\frac{1}{x}}{\frac{1}{x}+\frac{1}{\mu}}(x^{2}-\mu^{2})=\alpha_{f}^{\prime}(\mu-x)^{2}.\end{split}$
If $x=\mu$ then (5.9) becomes $0\geq 0$ and is still true. This completes the proof. ∎

### v2

@@ 6 Conclusion
(98) $\Omega_{n}=B_{R_{n}}(0)-\bigl(B_{\frac{1}{n}}(\theta_{a})\cup B_{\frac{1}{n}}(\theta_{b})\bigr)$
where $n>\frac{1}{\delta}$ and $R_{n}>n$ is large enough such that $B_{R_{n}}(0)$ contains both $B_{\delta}(\theta_{a})$ and $B_{\delta}(\theta_{b})$, $\int_{B_{R_{n}}^{c}(0)}\rho\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$, and $\int_{B_{R_{n}}^{c}(0)}{\rho^{*}}\mathrm{d}\theta<\frac{|B_{1/n}(0)|\epsilon}{2}$. Without loss of generality, we assume $\int_{\Omega_{n}}{\rho^{*}}\mathrm{d}\theta>\int_{\Omega_{n}}\rho\mathrm{d}\theta$, then we define
(99) $\Omega_{n}(r)=B_{R_{n}}(0)-\bigl(B_{r}(\theta_{a})\cup B_{1/n}(\theta_{b})\bigr)$
@@
(100) $\int_{{\Omega}_{n}(r)}{\rho^{*}}\mathrm{d}\theta$ $=1-\int_{B_{R_{n}}(0)^{c}}{\rho^{*}}\mathrm{d}\theta-\int_{B_{1/n}(\theta_{b})}{\rho^{*}}\mathrm{d}\theta$
$\leq 1-\int_{B_{1/n}(\theta_{b})}(\rho+\epsilon)\mathrm{d}\theta$
$=1-\int_{B_{1/n}(\theta_{b})}\rho\mathrm{d}\theta-\epsilon|B_{1/n}(0)|$
@@
Therefore, by continuity, there exists some $r_{n}\in(0,1/n)$, such that $\int_{{\Omega}_{n}(r_{n})}{\rho^{*}}\mathrm{d}\theta=\int_{{\Omega}_{n}(r_{n})}\rho\mathrm{d}\theta$. Using ${\Omega}_{n}(r_{n})$ to replace $\Omega_{n}$ completes the proof. We can also make $\Omega_{n}$ monotonically increasing by using the union $\cup_{k=1}^{n}\Omega_{k}$ to replace $\Omega_{n}$.
Proof A.2 (Proof of Lemma 4.6).
Assume that 4.2 holds with $\alpha>0$, namely
(101) $\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)^{2}\geq\alpha(\frac{1}{x}-\frac{1}{y})\bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr)$
holds for any $0<x<1<y$. Let $y$ in (101) approach $1$, we get
(102) $f^{\prime}(x)^{2}\geq\alpha\frac{f(x)}{x}\quad\textrm{ for }\quad x>0.$
(103) $f^{\prime\prime}(1)\geq\frac{\alpha}{2}.$
(104) $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x)\leq 2f^{\prime\prime}(1).$
(105) $f^{\prime}(1-\delta)\leq\frac{\delta^{\prime}}{2}f^{\prime\prime}(1).$
Also, the convexity of $f$ implies that $f^{\prime}(x)$ is non-decreasing and $\frac{f(x)}{1-x}$ is non-increasing. Lemma 4.6 requires that
(106) $\exists\alpha_{\delta}>0\quad\textrm{s.t.}\quad\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr)$
holds for all $x\in(1-\delta,1+\delta)$ and $y>0$. Instead we consider a relaxed setting where $x,y>0$, and $\{x,y\}\cap(1-\delta,1+\delta)\neq\emptyset$. Due to the symmetry of Eq. 106 with respect to $x$ and $y$, we assume without loss of generality that $x<y$ (when $x=y$, Eq. 106 holds with $0\geq 0$). Moreover, by 4.2, it suffices to consider the case where $1\not\in[x,y]$. In the following, we will prove that Eq. 106 holds under conditions $x<y$, $\{x,y\}\cap(1-\delta,1+\delta)\neq\emptyset$, and $1\not\in[x,y]$. We will then use the fact that $0\leq\delta\leq\delta^{\prime}$ to divide the analysis into three cases:
  1. Case 1
    . $1-2\delta^{\prime}<x<y<1+2\delta^{\prime}$,
  2. Case 2
    . $1<x<1+\delta^{\prime}<1+2\delta^{\prime}<y$,
  3. Case 3
    . $0<x<1-2\delta^{\prime}<1-\delta<y<1$.
For Case 1, (106) can be obtained with $\alpha_{\delta_{1}}=\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})^{2}}{4}$ from multiplying the following two inequalities
[…]
$\geq\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})^{2}}{2}\bigl(\frac{1}{x}-\frac{1}{y}\bigr),$
[…]
$=f^{\prime\prime}(1)(y-x)$
$\leq 2\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr).$
Here in the last inequality, we used (107).
For Case 2, (106) can be obtained with $\alpha_{\delta_{2}}=\frac{\delta^{\prime}f^{\prime\prime}(1)}{6}$ from multiplying the following two inequalities together with (104) leading to
(109) $f^{\prime}(y)-f^{\prime}(x)$ $=\int_{x}^{y}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1+\delta^{\prime}}^{1+2\delta^{\prime}}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\geq\frac{\delta^{\prime}f^{\prime\prime}(1)}{2}\bigl(\frac{1}{x}-\frac{1}{y}\bigr),$
$\frac{f(x)}{1-x}-\frac{f(y)}{1-y}$ $=\int_{1}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi-\int_{1}^{x}f^{\prime\prime}(\xi)\frac{x-\xi}{x-1}{\rm d}\xi\quad(\textrm{using $f(1)=f^{\prime}(1)=0$})$
(110) $=\frac{y-x}{(y-1)(x-1)}\int_{1}^{x}f^{\prime\prime}(\xi)(\xi-1){\rm d}\xi+\int_{x}^{y}f^{\prime\prime}(\xi)\frac{y-\xi}{y-1}{\rm d}\xi$
$\leq\delta^{\prime}f^{\prime\prime}(1)+f^{\prime}(y)-f^{\prime}(x)$
$\leq 3\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr).$
Here in the last inequality, we used (109).
For Case 3, (106) can be obtained with $\alpha_{\delta_{3}}=\frac{\alpha}{4}$. In fact, the right hand side of (106) can be bounded as
(111) $(\frac{1}{x}-\frac{1}{y})\bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr)\leq\frac{f(x)}{x},$
by taking $y=1$ on the left hand side, since both $\frac{1}{y}$ and $\frac{f(y)}{1-y}$ are decreasing. For the left hand side of Eq. 106, both $f^{\prime}(x)$ and $f^{\prime}(y)$ are negative. Using (104) and (105) leads to
(112) $-f^{\prime}(x)$ $=\int_{x}^{1}f^{\prime\prime}(\xi){\rm d}\xi\geq\int_{1-2\delta^{\prime}}^{1}f^{\prime\prime}(\xi)\mathrm{d}\xi\geq\delta^{\prime}f^{\prime\prime}(1)\geq-2f^{\prime}(1-\delta)\geq-2f^{\prime}(y).$
Using Eq. 112,Eq. 111, and Eq. 102 leads to
Finally, choosing $\delta>0$ as in Eqs. 104 and 105 and combining all three cases, we conclude that (106) holds with $\alpha_{\delta}=\min\{\frac{f^{\prime\prime}(1)(1-2\delta^{\prime})}{4},\frac{\delta^{\prime}f^{\prime\prime}(1)}{6},\frac{\alpha}{4}\}$.
Proof A.3 (Proof of Lemma 4.7).
In this lemma, we assume that there exists $0<\delta<1$ and $0<\alpha_{\delta}\leq\alpha$ such that
(113) $\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)^{2}\geq\alpha_{\delta}(\frac{1}{x}-\frac{1}{y})\bigl(\frac{f(x)}{1-x}-\frac{f(y)}{1-y}\bigr),$
holds for all $0<x<1<y$ (4.2), and for any $x\in(1-\delta,1+\delta)$ and $y>0$ (Lemma 4.6). We also have the constraint
(114) $\sum_{i=1}^{3}\rho_{i}=1,\quad\sum_{i=1}^{3}\frac{\rho_{i}}{x_{i}}=1,\quad\rho_{i}>0,\quad x_{i}>0.$
Our goal is to show
(115) $\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}\Bigl(f^{\prime}(x_{i})-f^{\prime}(x_{j})\Bigr)^{2}\geq\alpha_{f}\sum_{1\leq i,j\leq 3}\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl(\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr).$
We assume that $x_{1}\leq x_{2}\leq x_{3}$, then due to the constraint (114), we have either $x_{1}=x_{2}=x_{3}=1$ or $x_{1}<1<x_{3}$.
For the first scenario, (115) holds with any $\alpha_{f}$. For the second scenario, when $1-\delta<x_{2}<1+\delta$, the relationships between pairs $(x_{1},x_{2})$, $(x_{1},x_{3})$ and $(x_{2},x_{3})$ fall into the conditions described in (113). Thus, (115) holds with $\alpha_{f}=\alpha_{\delta}$. Now, we only need to verify the following cases:
  1. Case 1
    . $0<x_{1}<1<1+\delta\leq x_{2}\leq x_{3}$,
  2. Case 2
    . $0<x_{1}\leq x_{2}\leq 1-\delta<1<x_{3}$.
For Case 1, we will first show that
(116) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl(\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr)\geq\delta\rho_{2}\rho_{3}(\frac{1}{x_{2}}-\frac{1}{x_{3}})\Bigl(\frac{f(x_{2})}{1-x_{2}}-\frac{f(x_{3})}{1-x_{3}}\Bigr).$
In fact, the convexity of $f(x)$ and the condition $f(1)=0$ imply that $\frac{f(x)}{1-x}$ is decreasing. Consequently, $\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\geq\frac{f(x_{2})}{1-x_{2}}-\frac{f(x_{3})}{1-x_{3}}\geq 0$, hence Eq. 116 can be obtained if we can show $\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\geq\delta\rho_{2}(\frac{1}{x_{2}}-\frac{1}{x_{3}}).$ This can be achieved using the following estimate:
Here the equality follows from Eq. 114, and the last inequality uses the fact that $x_{2}\geq 1+\delta$ and $x_{3}>0$.
With (116), we can prove that (115) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. To do so, we denote
(117) $F_{i,j}=\rho_{i}\rho_{j}(\frac{1}{x_{i}}-\frac{1}{x_{j}})\Bigl(\frac{f(x_{i})}{1-x_{i}}-\frac{f(x_{j})}{1-x_{j}}\Bigr),$
then we have
[…]
which shows that (115) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$.
For Case 2, we will first show that
(118) $\rho_{1}\rho_{3}(\frac{1}{x_{1}}-\frac{1}{x_{3}})\Bigl(\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\Bigr)\geq\delta\rho_{2}\rho_{1}(\frac{1}{x_{1}}-\frac{1}{x_{2}})\Bigl(\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{2})}{1-x_{2}}\Bigr).$
In fact, the convexity of $f(x)$ and the condition $f(1)=0$ imply that $\frac{f(x)}{1-x}$ is decreasing. Consequently, $\frac{f(x_{1})}{1-x_{1}}-\frac{f(x_{3})}{1-x_{3}}\geq\frac{f(x_{2})}{1-x_{2}}-\frac{f(x_{3})}{1-x_{3}}\geq 0$. Combining with $\frac{1}{x_{1}}-\frac{1}{x_{3}}\geq\frac{1}{x_{1}}-\frac{1}{x_{2}}\geq 0$, this implies that (118) holds if $\rho_{3}\geq\delta\rho_{2}$. This can be derived by the following estimate:
Here, in the last inequality, we used the fact that $x_{2}\leq 1-\delta$. Next, we will prove that (115) holds with $\alpha_{f}=\frac{\alpha_{\delta}\delta}{1+\delta}$. Using the definition in (117), we have
[…]
Combining both Case 1 and Case 2, we complete the proof.
@@ 2. Case 2
(119) $f(y)\geq f(x)+(y-x)f^{\prime}(x),\quad\forall\ y>0,\ x>0.$
@@
(120) $\bigl(f^{\prime}(y)-f^{\prime}(x)\bigr)(y-x)\geq(y-x)f^{\prime}(y)-f(y)+f(x)\geq 0,\quad\forall\ y>0,\ x>0.$
Then, for $x>0$, using Eq. 120, we obtain
(121) $\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr)(\frac{1}{\mu}-\frac{1}{x})\geq$ $(\frac{1}{\mu}-\frac{1}{x})f^{\prime}(\frac{1}{\mu})-f(\frac{1}{\mu})+f(\frac{1}{x})\geq 0,$
@@
(122) $=$ $\mu(x-\mu)f^{\prime}(\frac{1}{\mu})-\mu^{2}x\bigl(f(\frac{1}{\mu})-f(\frac{1}{x})\bigr).$
The first inequality follows from $\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr)(\frac{1}{\mu}-\frac{1}{x})\geq 0$, and the second from Eq. 121. Taking $x={\rho^{*}}/\rho$ in (122) leads to
$\Bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{\rho}{{\rho^{*}}})\Bigr)\Bigl(\frac{{\rho^{*}}^{2}}{\rho^{2}}-\mu^{2}\Bigr)\geq\mu(\frac{{\rho^{*}}}{\rho}-\mu)f^{\prime}(\frac{1}{\mu})-\frac{\mu^{2}{\rho^{*}}}{\rho}\Bigl(f(\frac{1}{\mu})-f(\frac{\rho}{{\rho^{*}}})\Bigr),$
which leads to the desired inequality. This completes the proof.
Proof B.2 (Proof of Lemma 5.4).
Since $f\in C^{2}$ and $f^{\prime\prime}(1)>0$, there exists $0<\delta_{f}<\frac{1}{4}$, such that for any $x\in(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})$, we have $\frac{f^{\prime\prime}(1)}{2}\leq f^{\prime\prime}(x).$
For the left hand side of (84), for any $x,y$ satisfying $y>x>0$ and $\{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$, we have
(123) $f^{\prime}(\frac{1}{x})-f^{\prime}(\frac{1}{y})=\int_{1/y}^{1/x}f^{\prime\prime}(\xi){\rm d}\xi\geq\frac{f^{\prime\prime}(1)}{2}\int_{(\frac{1}{y},\frac{1}{x})\cap(\frac{1}{1+2\delta_{f}},\frac{1}{1-2\delta_{f}})}{\rm d}\xi$
@@
$\geq$ $\alpha_{\chi^{2},f}^{\prime}\frac{y-x}{y+x},$
where $\alpha_{\chi^{2},f}^{\prime}=\frac{f^{\prime\prime}(1)\delta_{f}}{2(1+2\delta_{f})(1+\delta_{f})}>0$. These cases cover all situations with $y>x>0,\ \{x,y\}\cap[1,1+\delta_{f}]\neq\emptyset$.
Then we consider (84) with $\mu\in[1,1+\delta_{f}]$. When $0<x<\mu$, by using (123), we have
(124) $\begin{split}\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr)(x^{2}-\mu^{2})\geq\alpha_{\chi^{2},f}^{\prime}\frac{\frac{1}{x}-\frac{1}{\mu}}{\frac{1}{x}+\frac{1}{\mu}}(\mu^{2}-x^{2})=\alpha_{\chi^{2},f}^{\prime}(\mu-x)^{2}.\end{split}$
When $\mu<x$, by using (123), we have
(125) $\begin{split}\bigl(f^{\prime}(\frac{1}{\mu})-f^{\prime}(\frac{1}{x})\bigr)(x^{2}-\mu^{2})\geq\alpha_{\chi^{2},f}^{\prime}\frac{\frac{1}{\mu}-\frac{1}{x}}{\frac{1}{x}+\frac{1}{\mu}}(x^{2}-\mu^{2})=\alpha_{\chi^{2},f}^{\prime}(\mu-x)^{2}.\end{split}$
If $x=\mu$ then (84) becomes $0\geq 0$ and is still true. This completes the proof.